Operations Report November 7th

SOL: 4

Crew Carmen Atauconcha

Generator— It is working perfectly. On.

Solar— Working perfectly

Diesel – 70%

Propane – 57psi

Water (trailer) –

Water (static) – 15cc from top

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter:

ATV’s Used: No

Oil Checked: No

Atv # Fuel Used Gals: Not used

Tires Status:

# Hours Used Day: No EVA’s today

Phobos used: No


Beginning charge:-

Ending charge:-

Currently charging: –

Deimos used: No


Beginning charge:-

Ending charge: –

Currently charging:-

HabCar used? No

Notes and Comments:

Today we had a meeting, where we decide how to reduce the usage of water. Also, we review the radios again and separate which ones are working. In some of them the sound is low. For tomorrow we are planning to organize and clean the tools and equipments that we have in the Hab.

Summary of operations activities:

Everyone is working in their researches. For tomorrow we are going to have an EVA session.

Questions and Concerns to Mission Support:

Hello mission support, I have a question about phobos and deimos, Which are the deimos, and which are the phobos? Please.

Copyright © The Mars Society. All rights reserved. | Main Site